19. Still using the same Demand and Supply curves that can be represented by:     Qd =...

19.

Still using the same Demand and Supply curves that can be represented by:     Qd = 90-2P; and Qs = 3P.   Suppose price is at $20, there will be ……..    as many as ……..units.

shortage; 10 units
shortage; 18 units
surplus; 10 units
surplus; 18 units
Answer #1

Solution:

Qd = 90-2P

Qs = 3P

At P = $20, quantity demanded or Qd = 90-2*20 = 50

quantity supplied Qs = 3P = 3*20 = 60

Since demand is less than supply, there is surplus of 10 units (60-50).